OpenStudy (anonymous):

can someone please show me how to get x by itself i got most of it done but i dont know how to do the rest. X^2+24X+90=0 X^2+24=-90 X^2+24=54 (X+12)^2=54 sqrt(X+12)^2=sqrt54 x+12=+-3sqrt6

OpenStudy (anonymous):

@terenzreignz

jimthompson5910 (jim_thompson5910):

subtract 12 from both sides to get \[\large x = -12 \pm 3\sqrt{6}\]

OpenStudy (anonymous):

is this my answer x=+-3-12sqrt6

jimthompson5910 (jim_thompson5910):

no it is \[\large x = -12 \pm 3\sqrt{6}\]

OpenStudy (anonymous):

how? subtracted 12 from both

jimthompson5910 (jim_thompson5910):

btw you have a typo on line 3, it should be x^2 + 24x + 144 = -90 + 144

jimthompson5910 (jim_thompson5910):

yeah you go from x+12=+-3sqrt6 to x = -12+-3sqrt6

jimthompson5910 (jim_thompson5910):

after you subtract 12 from both sides (you don't combine the terms since they aren't like terms)

OpenStudy (anonymous):

oh ok thanks i see what i did wrong

jimthompson5910 (jim_thompson5910):

ok great